OpenStudy (anonymous):

A garden table and a bench cost $750 combined. The cost of the garden table is two times the cost of the bench. What is the cost of the bench?

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

Let t = cost of garden table b = cost of bench The two combined are $750, so t + b = 750 The cost of the garden table is two times the cost of the bench, so t = 2b

jimthompson5910 (jim_thompson5910):

t + b = 750 2b + b = 750 ... plug in t = 2b 3b = 750 I'll let you finish up